# GRPO-MA: Multi-Answer Generation in GRPO for Stable and Efficient Chain-of-Thought Training

### v2

@@ 6 Conclusion
We present GRPO-MA, a simple yet theoretically grounded extension of GRPO that tackles three key challenges in training Chain-of-Thought models: unstable advantage estimation, gradient coupling between thoughts and answers, and sparse reward signals under limited sampling. By generating multiple answers per thought, GRPO-MA reduces the variance of advantage estimation, decouples the gradient between thoughts and answers, and densifies reward feedback. Our theoretical analysis further shows that increasing the number of answers per thought is a principled way to stabilize gradients, which is corroborated by experiments on math, code, and multimodal tasks. Together, these results demonstrate that GRPO-MA improves both the stability and efficiency of GRPO-based reinforcement learning.
Limitation Our study has several limitations. First, computational constraints prevent our experiments on larger-scale models. Second, our analysis relies on the simplifying assumption that thought values are independent, a condition that may not hold true in practice. Finally, the lack of a general-purpose reward model means that our testing is confined to tasks with verifiable rewards.
Appendix A Appendix
@@
  • More Related Work A.1
  • Full Analysis of Variance A.2
    • The Multivariate Delta Method A.2.1
    • Asymptotic Normality of the Estimated Value Vector A.2.2
    • Application to the Thought Advantage Function A.2.3
    • Application to the Answer Advantage Function A.2.4
    • Diagonality Analysis of Matrices A.2.5
  • Details in Task Settings A.3
    • Math A.3.1
    • Code A.3.2
    • Object Detection A.3.3
    • Affordance Prediction A.3.4
    • Trajectory Prediction A.3.5
    • Demand Prediction A.3.6
    • Ocr-based VQA A.3.7
    • Simulator-based Visual Manipulation A.3.8
  • Details in Training A.4
    • Training Hyperparameters A.4.1
    • SFT Details A.4.2
    • Geneartion Configure A.4.3
  • More Case Study and Visualization A.5
  • More Experimental Analysis A.6
    • Accuracy Reward Curve A.6.1
    • Richness of Reward Signal A.6.2
  • Usage of LLMs A.7
A.1 More Related Work: Applications of GRPO in Multimodal Domains
VLM-R1 Shen et al. (2025) applies a general GRPO pipeline to Vision-Language Models, enabling smaller models to achieve competitive performance on complex visual reasoning tasks. Vision-R1 Huang et al. (2025b) generates high-quality multimodal Chain-of-Thought data and uses Progressive Thinking Suppression Training (PTST) to prevent the model from creating overly long reasoning paths. Video-R1 Feng et al. (2025) introduces Temporal-GRPO (T-GRPO), a novel reward scheme that encourages the model to leverage temporal information in video sequences. ManipLVM-R1 Song et al. (2025) employs GRPO for robotic manipulation with new affordance-aware and trajectory matching reward functions to improve the localization of interactive parts and the physical plausibility of actions. Robot-R1 Kim et al. (2025) reframes robot learning as a multiple-choice question answering task, using GRPO to optimize the reasoning for embodied manipulation.
A.2 Full Analysis of Variance
This document provides a full derivation of the approximate variance for the Thought Advantage, $A(th_{i})$, as presented in the main paper. We first review the multivariate Delta Method, establish the asymptotic normality of our estimators via the Central Limit Theorem (CLT), and finally present the detailed application and gradient calculation.
A.2.1 The Multivariate Delta Method
@@
$\mathrm{Var}(f(\overrightarrow{V}_{M}))\approx\nabla f(\overrightarrow{\mu})^{T}\mathrm{Var}(\overrightarrow{V}_{M})\nabla f(\overrightarrow{\mu})$ (7)
where $\mathrm{Var}(\overrightarrow{V}_{M})$ is the actual covariance matrix of the estimator vector, which is related to the asymptotic covariance by $\mathrm{Var}(\overrightarrow{V}_{M})\approx\Sigma_{\text{asymptotic}}/M$.
A.2.2 Asymptotic Normality of the Estimated Value Vector
@@
A.2.3 Application to the Thought Advantage Function
@@
Finally, by the first-order multivariate Delta method, with $\mathrm{Var}(\overrightarrow{V})=\frac{1}{M}\Sigma_{\operatorname{diag}}$ (and $\Sigma_{\operatorname{diag}}=\operatorname{diag}(\sigma_{R_{1}}^{2},\dots,\sigma_{R_{K}}^{2})$),
$\mathrm{Var}[A(th_{i})]\approx\nabla_{V}f_{i}(\mu)^{\top}\mathrm{Var}(\overrightarrow{V})\,\nabla_{V}f_{i}(\mu)=\frac{1}{M\sigma_{\mu_{R}}^{2}}\sum_{k=1}^{K}\Big(\delta_{ik}-\frac{1}{K}-\frac{\tilde{\mu}_{i}\tilde{\mu}_{k}}{K-1}\Big)^{2}\sigma_{R_{k}}^{2}.$ (18)
A.2.4 Application to the Answer Advantage Function
@@
$\text{Var}\big[A(ans_{i,j})\big]\approx\nabla_{\mathbf{R}}g_{i,j}(\bm{\mu})^{\top}\,\text{diag}(\sigma_{R_{1}}^{2},\dots,\sigma_{R_{K}}^{2},\dots)\,\nabla_{\mathbf{R}}g_{i,j}(\bm{\mu}),$ (20)
where $g_{i,j}(\mathbf{R})=A(ans_{i,j})$ and $\bm{\mu}$ denotes the vector of reward means.
Evaluating the gradient at $\mathbf{R}=\bm{\mu}$ and grouping by thought, we obtain
@@
A.2.5 Diagonality Analysis of Matrices
@@
A.3 Details in Task Settings
A.3.1 Math
We conduct our experiments using problems from the DAPO Yu et al. (2025) training set and evaluate on the AIME2024 test set Maxwell-Jia (2024). The Math training set is constructed by randomly sampling 1,0001{,}000 problems from the DAPO training corpus. The model is trained for a single epoch on these 1,0001{,}000 training samples. We do not use a validation set; instead, we select the final model parameters saved at the end of training (the last checkpoint) for testing.
@@ A.3.1 Math
A.3.2 Code
We conduct our experiments using the Python-code portion of the SYNTHETIC-1 dataset PrimeIntellect (2024) and evaluate on the LiveBench code test set White et al. (2024). The Code training set is constructed by randomly sampling 1,0001{,}000 problems from the SYNTHETIC-1 Python-code corpus. The model is trained for a single epoch on these 1,0001{,}000 training samples. We do not use a validation set; instead, we select the final model parameters saved at the end of training (the last checkpoint) for testing.
@@ A.3.2 Code
A.3.3 Object Detection
We conduct our experiments using the Agibot World dataset contributors (2024). The data is partitioned into training, validation, and test sets based on specific task_ids from Agibot World dataset. Specifically, the training set is constructed from task_ids 424, 480, and 507, comprising a total of 3,000 images (randomly sampling). The validation and test sets are derived from task_id 582 and 1352, respectively. For all images, the ground-truth bounding boxes and corresponding object labels are annotated through a crowdsourcing process. The object detection model is trained for a single epoch on the 3,000-image training set.
@@ A.3.3 Object Detection
$\mathrm{IoU@\tau}=\frac{\sum_{i=1}^{N}\displaystyle\bm{1}(\text{IoU}(B_{pred}^{(i)},B_{gt}^{(i)})>\tau)}{N}$ (26)
where $N$ is the total number of samples in the test set, and $\bm{1}(\cdot)$ is the indicator function. To provide a comprehensive assessment, we report the performance across four different IoU thresholds: $\tau\in\{0.5,0.6,0.7,0.8\}$.
@@ A.3.4 Affordance Prediction
We primarily use the UMD Part Affordance Dataset Myers et al. (2015). The official training split of this dataset is used to construct our training and validation sets. Specifically, we use 3,000 images for training and a held-out portion of the original training split for validation. For evaluation, we use the official test split of the UMD dataset. To further assess the model’s generalization capabilities, we also use the entire AGD20K dataset Luo et al. (2022) as an additional, challenging test set.
@@
$\text{Success Rate}=\frac{\sum_{i=1}^{N}\displaystyle\bm{1}(M_{\text{gt}}^{(i)}(C_{\text{pred}}^{(i)})=1)}{N}$ (27)
where $N$ is the total number of samples in the test set, $C_{\text{pred}}^{(i)}$ is the predicted 2D coordinate $(x,y)$ for the $i$-th sample, and $M_{\text{gt}}^{(i)}$ is the corresponding ground-truth affordance mask. The notation $M_{\text{gt}}^{(i)}(C_{\text{pred}}^{(i)})$ represents the value of the mask at the predicted coordinate. $\bm{1}(\cdot)$ is the indicator function, which is 1 if the condition is true and 0 otherwise.
@@ A.3.5 Trajectory Prediction
We primarily use the trajectory subset of the BAAI ShareRobot dataset Ji et al. (2025). The original dataset is partitioned into training, validation, and test sets. Specifically, we use 3,000 images for training, a held-out portion of the training split for validation, and the test split for evaluation. The model is trained for a single epoch on the 3,000-image training set. After training, we perform model selection by evaluating checkpoints on the designated validation set. The checkpoint that achieves the highest reward value (as defined below) on the validation data is selected for the final evaluation. The performance of this selected model is then reported on the held-out test set.
We evaluate the model’s performance using multiple geometric similarity metrics, following the design in ManipVLM-R1 Song et al. (2025). These metrics measure how well the predicted trajectory matches the ground truth from different perspectives. Specifically, we use Discrete Fréchet Distance (DFD), Hausdorff Distance (HD), Root Mean Square Error (RMSE), and Endpoint Distance as evaluation criteria.
@@
The reward function consists of two complementary components: a format rewardand a accuracy reward. The format reward assigns a value of $1$ if and only if the output strictly adheres to this format; otherwise, it is set to 0. To measure how well the predicted trajectory $\hat{T}$ matches the ground-truth trajectory $T^{*}$, we adopt an accuracy reward following the design in ManipVLM-R1 Song et al. (2025). Specifically, the reward is defined as
@@ A.3.6 Demand Prediction
We construct the dataset for this task based on MO-DDN Wang et al. (2024), which requires robots to ground a natural demand instruction to objects in the environment. MO-DDN itself is built upon the HSSD scene dataset Khanna et al. (2024), together with a custom demand–object dataset. To build our data, we randomly sample a demand instruction and pair it with a scene containing a target object that satisfies the demand. We then crop and store the corresponding image, resulting in instruction–image pairs.
@@
$\text{Success Rate}=\frac{\sum_{i=1}^{N}\displaystyle\bm{1}(M_{\text{gt}}^{(i)}(C_{\text{pred}}^{(i)})=1)}{N},$ (29)
where $N$ is the number of samples in the test set, $C_{\text{pred}}^{(i)}$ denotes the predicted 2D coordinate $(x,y)$ for the $i$-th sample, and $M_{\text{gt}}^{(i)}$ is the ground-truth binary mask of the demanded object. The notation $M_{\text{gt}}^{(i)}(C_{\text{pred}}^{(i)})$ indicates the mask value at the predicted location. $\bm{1}(\cdot)$ is the indicator function that equals $1$ if the condition holds and 0 otherwise.
@@ A.3.7 OCR-based VQA
We construct the dataset by combining three OCR-based VQA benchmarks: Document VQA Tito et al. (2021), Infographics VQA Tito et al. (2021), and Scene Text VQA Biten et al. (2019). Document VQA focuses on answering questions asked over document images, which may contain printed, typewritten, and handwritten content (e.g., letters, memos, reports). The answers are typically text spans taken verbatim from the document. Infographics VQA considers questions over infographic images containing charts, diagrams, or other structured visual data, where answers are not always explicitly extracted text but can include inferred information. Scene Text VQA consists of natural scene images with embedded text (e.g., storefronts, street signs). The model must jointly leverage OCR reading and visual understanding to answer the questions.
@@
A.3.8 Simulator-based Visual Manipulation
@@
We construct the dataset and evaluation splits based on the PartNet Mobility dataset Xiang et al. (2020) and the ManipLLM experimental setup (A crucial point is that we have followed their setting by using suckers as the end effectors for the robotic arms.). For training, we adopt the same 20 training categories as ManipLLM, consisting of 1,043 object instances. Training scenes are generated following the SAPIEN simulator Xiang et al. (2020) setup and ManipLLM scene configurations. For testing, we use the open-sourced ManipLLM test set, which contains approximately 1,830 successful test samples spanning both Seen and Unseen objects. To better evaluate model generalization to novel viewpoints, we further construct a camera-perturbed test set by modifying each test sample: the camera orientation vector $[0,0,0]$ is replaced by $[x,y,z]$ where each of $x,y,z$ is sampled uniformly from the signed interval $\pm[0.2,0.6]$. This perturbation preserves other scene properties while intentionally stressing viewpoint robustness. In order to simplify control and isolate contact selection, the sucker approach direction in all experiments is fixed to be the surface normal at the chosen manipulation point $(x,y)$.
@@
$\text{SuccessRate}=\frac{1}{N}\sum_{i=1}^{N}\displaystyle\bm{1}\{\text{trial}_{i}\text{ is successful according to ManipLLM's displacement criterion}\},$
where $\bm{1}\{\cdot\}$ is the indicator function. We report Success Rate on the camera-perturbed test sets, and further provide breakdowns by Seen vs. Unseen objects.
@@
A.4 Details in Training
A.4.1 Training Hyperparameters

### v3

@@ 6 Conclusion
Under a simplified tree-branching setting where multiple answers are sampled per thought, we provide a variance-based analysis of thought-level advantage estimation in GRPO-style training. Our analysis reveals a fundamental asymmetry: increasing the number of answers per thought $M$ monotonically decreases the variance toward zero, whereas increasing the number of sampled thoughts $K$ leaves a strictly positive variance floor. Accordingly, under GRPO-style estimators without value models and absent additional mechanisms for controlling thought generation, multi-answer branching emerges as a effective and potentially necessary means for achieving stable and accurate advantage estimation. Empirically, GRPO-MA, a simple instantiation of this principle, consistently improves performance and optimization stability across different tasks, model configurations.
Impact Statement
This paper presents work whose goal is to advance the field of Large Language Models. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Appendix Outline
@@
  • More Related Work B
  • Full Analysis of Variance C
    • The Multivariate Delta Method C.1
    • Asymptotic Normality of the Estimated Value Vector C.2
    • Application to the Thought Advantage Function C.3
    • Application to the Answer Advantage Function C.4
    • Diagonality Analysis of Matrices C.5
  • Details in Task Settings D
    • Math D.1
    • Code D.2
    • Object Detection D.3
    • Affordance Prediction D.4
    • Trajectory Prediction D.5
    • Demand Prediction D.6
    • Ocr-based VQA D.7
    • Simulator-based Visual Manipulation D.8
  • Details in Training E
    • Training Hyperparameters E.1
    • SFT Details E.2
    • Geneartion Configure E.3
  • Case Study and Visualization F
  • More Experimental Analysis G
    • Inconsistency Analysis G.1
    • Accuracy Reward Curve G.2
    • Richness of Reward Signal G.3
    • Wall-clock Time G.4
    • Grad-norm and GSS Curve G.5
    • More Results on Code and Math G.6
  • Usage of LLMs H
Appendix B More Related Work: Applications of GRPO in Multimodal Domains
VLM-R1 (Shen et al., 2025) applies a general GRPO pipeline to Vision-Language Models, enabling smaller models to achieve competitive performance on complex visual reasoning tasks. Vision-R1 (Huang et al., 2025b) generates high-quality multimodal Chain-of-Thought data and uses Progressive Thinking Suppression Training (PTST) to prevent the model from creating overly long reasoning paths. Video-R1 (Feng et al., 2025) introduces Temporal-GRPO (T-GRPO), a novel reward scheme that encourages the model to leverage temporal information in video sequences. ManipLVM-R1 (Song et al., 2025) employs GRPO for robotic manipulation with new affordance-aware and trajectory matching reward functions to improve the localization of interactive parts and the physical plausibility of actions. Robot-R1 (Kim et al., 2025) reframes robot learning as a multiple-choice question answering task, using GRPO to optimize the reasoning for embodied manipulation.
Appendix C Full Analysis of Variance
This section provides a full derivation of the approximate variance for the Thought Advantage, $A(th_{i})$, as presented in the main paper. We first review the multivariate Delta Method, establish the asymptotic normality of our estimators via the Central Limit Theorem (CLT), and finally present the detailed application and gradient calculation.
C.1 The Multivariate Delta Method
@@
$\operatorname{Var}(f(\overrightarrow{V}_{M}))\approx\nabla f(\overrightarrow{\mu})^{T}\operatorname{Var}(\overrightarrow{V}_{M})\nabla f(\overrightarrow{\mu})$ (6)
where $\operatorname{Var}(\overrightarrow{V}_{M})$ is the actual covariance matrix of the estimator vector, which is related to the asymptotic covariance by $\operatorname{Var}(\overrightarrow{V}_{M})\approx\Sigma_{\text{asymptotic}}/M$.
C.2 Asymptotic Normality of the Estimated Value Vector
@@
C.3 Application to the Thought Advantage Function
@@
Finally, by the first-order multivariate Delta method, with $\operatorname{Var}(\overrightarrow{V})=\frac{1}{M}\Sigma_{\operatorname{diag}}$ (and $\Sigma_{\operatorname{diag}}=\operatorname{diag}(\sigma_{R_{1}}^{2},\dots,\sigma_{R_{K}}^{2})$),
$\operatorname{Var}[A(th_{i})]\approx\nabla_{V}f_{i}(\mu)^{\top}\operatorname{Var}(\overrightarrow{V})\,\nabla_{V}f_{i}(\mu)=\frac{1}{M\sigma_{\mu_{R}}^{2}}\sum_{k=1}^{K}\Big(\delta_{ik}-\frac{1}{K}-\frac{\tilde{\mu}_{i}\tilde{\mu}_{k}}{K-1}\Big)^{2}\sigma_{R_{k}}^{2}.$ (17)
C.4 Application to the Answer Advantage Function
@@
$\text{Var}\big[A(ans_{i,j})\big]\approx\nabla_{\mathbf{R}}g_{i,j}(\boldsymbol{\mu})^{\top}\,\text{diag}(\sigma_{R_{1}}^{2},\dots,\sigma_{R_{K}}^{2},\dots)\,\nabla_{\mathbf{R}}g_{i,j}(\boldsymbol{\mu}),$ (19)
where $g_{i,j}(\mathbf{R})=A(ans_{i,j})$ and $\boldsymbol{\mu}$ denotes the vector of reward means.
Evaluating the gradient at $\mathbf{R}=\boldsymbol{\mu}$ and grouping by thought, we obtain
@@
C.5 Diagonality Analysis of Matrices
@@
Appendix D Details in Task Settings
D.1 Math
We conduct our experiments using problems from the DAPO (Yu et al., 2025) training set and evaluate on the AIME2024 test set (Maxwell-Jia, 2024). The Math training set is constructed by randomly sampling 1,0001{,}000 problems from the DAPO training corpus. The model is trained for a single epoch on these 1,0001{,}000 training samples. We do not use a validation set; instead, we select the final model parameters saved at the end of training (the last checkpoint) for testing.
@@ D.1 Math
D.2 Code
We conduct our experiments using the Python-code portion of the SYNTHETIC-1 dataset (PrimeIntellect, 2024) and evaluate on the LiveBench code test set (White et al., 2024). The Code training set is constructed by randomly sampling 1,0001{,}000 problems from the SYNTHETIC-1 Python-code corpus. The model is trained for a single epoch on these 1,0001{,}000 training samples. We do not use a validation set; instead, we select the final model parameters saved at the end of training (the last checkpoint) for testing.
@@ D.2 Code
D.3 Object Detection
We conduct our experiments using the Agibot World dataset (contributors, 2024). The data is partitioned into training, validation, and test sets based on specific task_ids from Agibot World dataset. Specifically, the training set is constructed from task_ids 424, 480, and 507, comprising a total of 3,000 images (randomly sampling). The validation and test sets are derived from task_id 582 and 1352, respectively. For all images, the ground-truth bounding boxes and corresponding object labels are annotated through a crowdsourcing process. The object detection model is trained for a single epoch on the 3,000-image training set.
@@ D.3 Object Detection
$\mathrm{IoU@\tau}=\frac{\sum_{i=1}^{N}\mathbb{I}(\text{IoU}(B_{pred}^{(i)},B_{gt}^{(i)})>\tau)}{N}$ (25)
where $N$ is the total number of samples in the test set, and $\mathbb{I}(\cdot)$ is the indicator function. To provide a comprehensive assessment, we report the performance across four different IoU thresholds: $\tau\in\{0.5,0.6,0.7,0.8\}$.
@@ D.4 Affordance Prediction
We primarily use the UMD Part Affordance Dataset (Myers et al., 2015). The official training split of this dataset is used to construct our training and validation sets. Specifically, we use 3,000 images for training and a held-out portion of the original training split for validation. For evaluation, we use the official test split of the UMD dataset. To further assess the model’s generalization capabilities, we also use the entire AGD20K dataset (Luo et al., 2022) as an additional, challenging test set.
@@
$\text{Success Rate}=\frac{\sum_{i=1}^{N}\mathbb{I}(M_{\text{gt}}^{(i)}(C_{\text{pred}}^{(i)})=1)}{N}$ (26)
where $N$ is the total number of samples in the test set, $C_{\text{pred}}^{(i)}$ is the predicted 2D coordinate $(x,y)$ for the $i$-th sample, and $M_{\text{gt}}^{(i)}$ is the corresponding ground-truth affordance mask. The notation $M_{\text{gt}}^{(i)}(C_{\text{pred}}^{(i)})$ represents the value of the mask at the predicted coordinate. $\mathbb{I}(\cdot)$ is the indicator function, which is 1 if the condition is true and 0 otherwise.
@@ D.5 Trajectory Prediction
We primarily use the trajectory subset of the BAAI ShareRobot dataset (Ji et al., 2025). The original dataset is partitioned into training, validation, and test sets. Specifically, we use 3,000 images for training, a held-out portion of the training split for validation, and the test split for evaluation. The model is trained for a single epoch on the 3,000-image training set. After training, we perform model selection by evaluating checkpoints on the designated validation set. The checkpoint that achieves the highest reward value (as defined below) on the validation data is selected for the final evaluation. The performance of this selected model is then reported on the held-out test set.
We evaluate the model’s performance using multiple geometric similarity metrics, following the design in ManipVLM-R1 (Song et al., 2025). These metrics measure how well the predicted trajectory matches the ground truth from different perspectives. Specifically, we use Discrete Fréchet Distance (DFD), Hausdorff Distance (HD), Root Mean Square Error (RMSE), and Endpoint Distance as evaluation criteria.
@@
The reward function consists of two complementary components: a format rewardand a accuracy reward. The format reward assigns a value of $1$ if and only if the output strictly adheres to this format; otherwise, it is set to 0. To measure how well the predicted trajectory $\hat{T}$ matches the ground-truth trajectory $T^{*}$, we adopt an accuracy reward following the design in ManipVLM-R1 (Song et al., 2025). Specifically, the reward is defined as
@@ D.6 Demand Prediction
We construct the dataset for this task based on MO-DDN (Wang et al., 2024), which requires robots to ground a natural demand instruction to objects in the environment. MO-DDN itself is built upon the HSSD scene dataset (Khanna et al., 2024), together with a custom demand–object dataset. To build our data, we randomly sample a demand instruction and pair it with a scene containing a target object that satisfies the demand. We then crop and store the corresponding image, resulting in instruction–image pairs.
@@
$\text{Success Rate}=\frac{\sum_{i=1}^{N}\mathbb{I}(M_{\text{gt}}^{(i)}(C_{\text{pred}}^{(i)})=1)}{N},$ (28)
where $N$ is the number of samples in the test set, $C_{\text{pred}}^{(i)}$ denotes the predicted 2D coordinate $(x,y)$ for the $i$-th sample, and $M_{\text{gt}}^{(i)}$ is the ground-truth binary mask of the demanded object. The notation $M_{\text{gt}}^{(i)}(C_{\text{pred}}^{(i)})$ indicates the mask value at the predicted location. $\mathbb{I}(\cdot)$ is the indicator function that equals $1$ if the condition holds and 0 otherwise.
@@ D.7 OCR-based VQA
We construct the dataset by combining three OCR-based VQA benchmarks: Document VQA (Tito et al., 2021), Infographics VQA (Tito et al., 2021), and Scene Text VQA (Biten et al., 2019). Document VQA focuses on answering questions asked over document images, which may contain printed, typewritten, and handwritten content (e.g., letters, memos, reports). The answers are typically text spans taken verbatim from the document. Infographics VQA considers questions over infographic images containing charts, diagrams, or other structured visual data, where answers are not always explicitly extracted text but can include inferred information. Scene Text VQA consists of natural scene images with embedded text (e.g., storefronts, street signs). The model must jointly leverage OCR reading and visual understanding to answer the questions.
@@
D.8 Simulator-based Visual Manipulation
@@
We construct the dataset and evaluation splits based on the PartNet Mobility dataset (Xiang et al., 2020) and the ManipLLM experimental setup (A crucial point is that we have followed their setting by using suckers as the end effectors for the robotic arms.). For training, we adopt the same 20 training categories as ManipLLM, consisting of 1,043 object instances. Training scenes are generated following the SAPIEN simulator (Xiang et al., 2020) setup and ManipLLM scene configurations. For testing, we use the open-sourced ManipLLM test set, which contains approximately 1,830 successful test samples spanning both Seen and Unseen objects. To better evaluate model generalization to novel viewpoints, we further construct a camera-perturbed test set by modifying each test sample: the camera orientation vector $[0,0,0]$ is replaced by $[x,y,z]$ where each of $x,y,z$ is sampled uniformly from the signed interval $\pm[0.2,0.6]$. This perturbation preserves other scene properties while intentionally stressing viewpoint robustness. In order to simplify control and isolate contact selection, the sucker approach direction in all experiments is fixed to be the surface normal at the chosen manipulation point $(x,y)$.
@@
$\text{SuccessRate}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\{\text{trial}_{i}\text{ is successful according to ManipLLM's displacement criterion}\},$
where $\mathbb{I}\{\cdot\}$ is the indicator function. We report Success Rate on the camera-perturbed test sets, and further provide breakdowns by Seen vs. Unseen objects.
@@
Appendix E Details in Training
E.1 Training Hyperparameters